OpenStudy (anonymous):

Show an equation and a solution for the problem. How much water must be added to 12 L (liters) of a 40% solution of alcohol to obtain a 30% solution?

OpenStudy (anonymous):

(0x)+12*.4=(12+x).3 I am not sure

OpenStudy (sriram):

In 12L there is 40%alcohol hence amount of alcohol=12*40/100 =4.8L

OpenStudy (sriram):

4.8L=30% hence 4.8= total volume*30/100 total volume=16L

OpenStudy (sriram):

therefore water added=4L

OpenStudy (anonymous):

thank you very much for the help I am not great at math you are truley life savers : )